### v1

@@ Abstract
Large language models (LLMs) are increasingly used as human simulators, both for evaluating conversational systems and for generating fine-tuning data. However, naive “act-as-a-user" prompting often yields verbose, unrealistic utterances, underscoring the need for principled evaluation of so-called user proxy agents. We present MirrorBench, a reproducible, extensible benchmarking framework that evaluates user proxies solely on their ability to produce human-like user utterances across diverse conversational tasks, explicitly decoupled from downstream task success. MirrorBench features a modular execution engine with typed interfaces, metadata-driven registries, multi-backend support, caching, and robust observability. The system supports pluggable user proxies, datasets, tasks, and metrics, enabling researchers to evaluate arbitrary simulators under a uniform, variance-aware harness. We include three lexical-diversity metrics (MATTR, Yule’s $K$, and HD-D) and three LLM-judge–based metrics (GTEval, Pairwise Indistinguishability, and Rubric-and-Reason). Across four open datasets, MirrorBench yields variance-aware results and reveals systematic gaps between user proxies and real human users. The framework is open source and includes a simple command-line interface for running experiments, managing configurations and caching, and generating reports. The framework can be accessed at https://github.com/SAP/mirrorbench
footnotetext: 1Anonymous Institution, Anonymous City, Anonymous Region, Anonymous Country. Correspondence to: Anonymous Author <anon.email@domain.com>.
Preliminary work. Under review by the Machine Learning and Systems (MLSys) Conference. Do not distribute.
1 Introduction & Related Work
The idea of simulating users to train and evaluate dialogue agents has a long history in NLP Balog and Zhai (2023). Early user simulators were typically goal-driven or rule-based, following predefined agendas of dialogue acts to mimic human behavior Schatzmann and Young (2009). Recent work has moved towards large language model (LLM) based user proxies that generate more realistic, open-ended interactions. For example, Wang et al. (2025) introduced a User Simulator with Implicit Profiles (USP) that infers latent user traits to produce personalized, authentic dialogues, addressing issues of role confusion and limited diversity in naive LLM role-play. Likewise, DialogueForge Zhu et al. (2025) bootstraps AI-to-AI conversations from real chat logs as seed prompts, leveraging powerful models (e.g. GPT-4) as the user to yield multi-turn dialogues. It demonstrates that while larger LLMs produce the most human-like exchanges, fine-tuned smaller models can approximate realism with customization, though maintaining long-range coherence remains a challenge. Collectively, these simulators underscore the potential of LLM-driven user agents for producing high-fidelity dialogues.
Parallel to improvements in user simulation, researchers have begun using LLMs as judges to evaluate dialogue quality. These “LLM-as-a-judge" evaluators prompt a strong model to rate or compare chatbot responses, bypassing the need for reference answers or extensive human annotation Liu et al. (2023a), Li et al. (2023). With careful prompt design (e.g. asking the model to explain its criteria via chain-of-thought reasoning), such LLM-based evaluation metrics can flexibly assess custom quality criteria Shankar et al. (2024). Notably, LLM evaluators have enabled large-scale dialogue assessments by providing fast, consistent scoring of responses. However, existing LLM-as-judge systems mainly focus on rating the assistant’s outputs rather than evaluating the simulators themselves.
@@ 1 Introduction & Related Work
Deploying and improving conversational AI hinges on evaluation under realistic user interactions. Expert human annotation, while authoritative, is costly and difficult to scale for systematic evaluation. At the same time, the instruction-tuning of models for production systems demands large volumes of authentic user–assistant exchanges. To meet both needs, practitioners increasingly rely on user proxies, namely LLMs prompted to emulate specified human personas, to synthesize user utterances for automated testing and post-training at scale. For example, recent work shows that LLM simulator can be used to evaluate an AI assistant over hundreds of interactive conversations Lu et al. (2025), Ahmad et al. (2025), Hathidara et al. (2025).
Yet the naive prompting of a single LLM to act as user simulator often yields verbose, overly cooperative behavior that diverges from real user speech patterns Zhou et al. (2024). Recent studies further note that assistant LLMs can lose coherence in extended dialogues, in part because they fail to capture or elicit the subtleties of human behavior Laban et al. (2025). Scalable, practical evaluation of conversational AI therefore depends on realistic user simulators, which must themselves be rigorously evaluated for how closely they resemble human users across diverse scenarios. This leads to the central question: How can we rigorously measure the human-likeness or user-likeness of a user proxy, independently of downstream task success? Here, we define a user as a human interlocutor participating in a specific conversational or task-oriented dialogue.
In this work, we introduce MirrorBench, a reproducible and extensible framework for benchmarking user-proxy agents on human-likeness metrics using multiple human reference datasets across diverse conversational settings. MirrorBench evaluates proxies solely on human-likeness and deliberately separates this from task completion. The framework comprises strongly typed domain models and metadata-rich registries, a compatibility-aware planner that validates dataset and metric compatibility while producing replayable manifests for experiment reproducibility, and a flexible execution engine supporting multi-turn interactions with synchronous, asynchronous, and distributed backends such as Ray Moritz et al. (2018). It also includes a caching layer for LLM outputs and dataset statistics, and a command-line interface (CLI) for validating, dry-running, executing, and reporting evaluation experiments. Together, these components form a modular harness in which proxies, datasets, tasks, and metrics are pluggable yet rigorously validated.
MirrorBench is designed for extensibility and controlled variation. Researchers can register custom LLM clients, user-proxy adapters, datasets, evaluation metrics, task drivers, and execution back-ends through a unified API. Component capabilities and compatibility constraints are declared in metadata (while registering) and validated by the planner, allowing modules to be swapped or extended without modifying the core. This separation of concerns enables fair, apples-to-apples comparisons across proxies and judges under consistent orchestration and logging, and ensures seamless integration into the evaluation pipeline.
@@
To evaluate the realism of user–proxy agents against real-world dialogues, MirrorBench packages four open-source corpora preprocessed for user-proxy evaluation: QULAC Aliannejadi et al. (2019), ClariQ Aliannejadi et al. (2021), OASST1 Köpf et al. (2023), and ChatbotArena Zheng et al. (2023). To quantify alignment in lexical diversity with human users, we implement Moving-Average Type–Token Ratio (MATTR) Covington and McFall (2010), Yule’s $K$ Tanaka-Ishii and Aihara (2015), and a Hypergeometric Distribution Diversity (HD-D) Carmona et al. (2019). To assess behavioral indistinguishability, we employ a family of LLM-judge metrics, including GTEval Zhu et al. (2025), Pairwise Indistinguishability (PI) Zheng et al. (2023), and Rubric-and-Reason (RNR) Liu et al. (2023b). Taken together, these metrics quantify the realism of the user side of the dialogue, explicitly decoupled from assistant capabilities and task success.
MirrorBench is not merely a metric suite; it is a full-fledged evaluation harness. It validates inputs against strongly typed schemas, persists every turn, score, and judge invocation to a run database, and caches LLM outputs for reproducibility and cost control. The engine logs fine-grained telemetry (latency, token counts, cost) at each turn and supports synchronous, asynchronous, and distributed evaluation. A compatibility-aware planner expands grids over proxies, datasets, metrics, and random seeds, and emits a replayable manifest that precisely reproduces an evaluation. The command-line interface supports quick dry-runs for validation, long-running jobs for large-scale experiments, and artifact inspection for post-hoc analysis. Structured logging via OpenTelemetry OpenTelemetry (2025) captures detailed traces for debugging and performance profiling. Together, these capabilities make MirrorBench both a rigorous benchmark of human-likeness and a scalable harness for stress-testing user proxies in a conversational setting.
In the next sections, we formalize the evaluation problem (§2), introduce the MirrorBench architecture and execution flow (§3), define metrics (§4) and datasets (§5), and detail the experimental setup and results (§6).
2 Problem Definition
@@
To compare a user–proxy agent’s behavior against human references, we synthesize proxy–assistant dialogues by rolling out the user–proxy $\theta_{u}$ against an assistant model $\theta_{a}$. For each reference dialogue $d^{\mathrm{ref}}_{j}$ with user goal $g_{j}$, we condition $\theta_{u}$ on $g_{j}$ and the accumulated dialogue history to produce user turns $\hat{u}_{j,t}$, while $\theta_{a}$ produces contextual assistant replies $\hat{a}_{j,t}$. This yields a synthetic transcript
@@
3 System Architecture
We present MirrorBench, a domain-agnostic framework for evaluating the human-likeness of user-proxy agents. Its architecture comprises six stacked layers. This section details the system design and end-to-end execution flow.
3.1 Overview
Figure 1: MirrorBench Architecture: Six-layer stack from low-level execution backends & persistence up through the core engine, plugin components, CLI & reporting, and task drivers. Top layers are user-facing; bottom layers are low-level infrastructure abstractions.
MirrorBench is a six-layer stack that cleanly separates infrastructure from evaluation logic (Fig. 1). Layers build upward, lower layers provide execution and data services; upper layers host pluggable components, task logic, and user interfaces. Full component details appear in the appendix A; here, we summarize each layer.
Execution Backends & Persistence (Layer 1): This base layer runs evaluation units and manages durable state. The execution substrate supports synchronous and asynchronous runners (with an extension point for Ray Moritz et al. (2018)) and a persistence plane (SQLite + files) that stores runs, episodes, metrics, telemetry, and caches.
Core Engine (Layer 2): This layer defines typed data models for all execution artifacts (e.g., messages, metric and dataset parameters), parses run configuration, and emits run manifests to ensure reproducibility. A metadata-aware registry and validators enforce compatibility constraints prior to execution, preventing silent integration errors.
Orchestration (Layer 3): The planner expands a job configuration into compatible (proxy, dataset, metric, seed) evaluation units; the run controller schedules them on the chosen backend (sync/async/ray), handles retries/timeouts, and records lineage. A cache layer deduplicates judge/model calls, while observability provides structured logging and optional tracing/metrics.
Plugin Components (Layer 4): All domain logic is modular: model clients (provider SDK wrappers with unified I/O and telemetry), user-proxy adapters (LLMs or agents behind interface), datasets (normalized episode loaders with manifests), and metrics (lexical and LLM-judge families with declared dependencies).
Task Drivers (Layer 5): Drivers encode the interaction protocol between user-proxy and assistant for a given task. We ship single-turn and multi-turn “mirror conversation” drivers that emit standardized artifacts consumable by metrics.
API & CLI (Layer 6): A thin programmatic API facade exposes plan/run/report operations; the CLI mirrors these (plan, dryrun, run, report, cache). Outputs include JSON summaries with aggregated metric results and confidence intervals, suitable for post-hoc analysis.
3.2 Execution Flow
Figure 2: MirrorBench execution flow. The framework decomposes an evaluation job into units $\{U_{1},\ldots,U_{p}\}$; each unit $U_{i}$ iterates over episodes $\{e_{1},\ldots,e_{n}\}$ produced by the dataset adapter and executed via task drivers. Metrics are computed per episode and aggregated within each unit with confidence intervals.
This section details the end-to-end execution that instantiates the pipeline $\Psi_{T}$ (as described in §2). Figure 2 shows the path from CLI invocation through planning, orchestration, and nested execution to persisted results, compiling high-level specifications into reproducible evaluation artifacts.
CLI Invocation and Planning. An evaluation job is declared via a configuration file (appendix E.2) that specifies user proxies $\mathcal{P}=\{\theta_{u_{1}},\ldots,\theta_{u_{q}}\}$, datasets $\mathcal{D}=\{D_{1},\ldots,D_{d}\}$, the metric set $M=\{m_{1},\ldots,m_{k}\}$, a fixed assistant model $\theta_{a}$, and run parameters (backend, observability, maximum concurrency, and a seed set $S=\{s_{1},\ldots,s_{r}\}$). The seed set enables repeated trials of the same combination under distinct random initializations, supporting variance estimation and confidence intervals.
Here $\operatorname{Compat}(\cdot)$ returns true iff the components can be jointly executed under their declared capabilities and requirements (e.g. metric prerequisites, model/back-end constraints). Each unit $U_{i}=(\theta_{u_{x}},D_{y},m_{z},s_{w})$ is executed and evaluated independently, with models, prompts, and driver assignments fixed at plan time. The planner creates & persists a replayable manifest file (appendix E.2) capturing all requisite metadata, and exposes a dryrun checkpoint for inspection prior to committing computational resources.
Orchestration and Backend Selection. The Runner instantiates a Run Controller that manages execution state, telemetry, and result persistence. The controller initializes a run-specific SQLite database to store task driver outputs and metric values. Based on the run configuration, the system selects an available backend (sync, async, or distributed) and dispatches the list of units $U$ for execution under the chosen strategy. For each unit $U_{i}\in U$, the backend invokes a unit executor that loads the dataset, instantiates the proxy, initializes the metrics, and configures the task driver.
Unit Execution and Episode Extraction. For each unit $U_{i}=(\theta_{u_{x}},D_{y},m_{z},s_{w})$, the unit executor loads dataset $D_{y}$ and enumerates its constituent episodes. Each episode $e_{j}$ is an immutable specification of a single reference dialogue (§2), with dataset layout and fields given by
$D_{y}$ $=\;\{\,e_{1},e_{2},\ldots,e_{n}\,\},$
$e_{j}$ $=\;\bigl(h_{j}^{\mathrm{ref}},\,g_{j},\,\mathrm{meta}_{j}\bigr),$
Conversation Synthesis via Task Drivers. For each episode $e_{j}$, the task driver orchestrates the synthesis of a proxy-assistant interaction by alternating invocations of the user proxy $\theta_{u_{x}}$ (abbrev. $\theta_{u}$) and the assistant model $\theta_{a}$, matching the reference dialogue length $L_{j}=|h_{j}^{\mathrm{ref}}|$. Let $\hat{h}_{j,t}$ denote the synthetic history after $t$ turns, with $\hat{h}_{j,0}=\langle\,\rangle$. The synthetic rollout for episode $e_{j}$ is:
$\textbf{Initialize:}\quad\hat{h}_{j,0}\leftarrow\langle\,\rangle$
$\textbf{For }t=1,\ldots,L_{j}\textbf{:}$
$r_{j}\;=\;\bigl(\,\hat{h}_{j,L_{j}},\;e_{j},\;\mathrm{telemetry}_{j}\bigr),$
Metric Computation. After producing an episode artifact $r_{j}$, metric $m_{z}$ evaluates $r_{j}$ to yield a scalar value $v_{j,m_{z}}$. Depending on the metric type (e.g., lexical, or LLM-judge), the computation may involve reference comparison, pairwise scoring, or external model inference. Formally,
$v_{j,m_{z}}\;=\;m_{z}\!\bigl(r_{j}\bigr),$
Result Aggregation. After processing all episodes $\{e_{1},\ldots,e_{n}\}$ within unit $U_{i}$, the executor aggregates the metric outputs into summary statistics. For metric $m_{z}$, let $V_{U_{i},m_{z}}=\{v_{1,m_{z}},v_{2,m_{z}},\ldots,v_{n,m_{z}}\}$ denote the set of episode-level scores. The unit-level aggregate includes the mean, standard deviation, and $95\%$ confidence interval:
$\bar{v}_{U_{i},m_{z}}\;$ $=\;\frac{1}{n}\sum_{j=1}^{n}v_{j,m_{z}},$
$s_{U_{i},m_{z}}\;$ $=\;\sqrt{\frac{1}{n}\sum_{j=1}^{n}(v_{j,m_{z}}-\bar{v}_{U_{i},m_{z}})^{2}},$
Iteration and Reporting. The backend executes each unit $U_{i}=(\theta_{u_{x}},D_{y},m_{z},s_{w})$ in $U=\{U_{1},\ldots,U_{p}\}$. Upon completion, the controller updates the last-run pointer for convenient access and consolidates a unified RunSummary that collates, for every $m_{z}$ and unit $U_{i}$, the tuple $\bigl(\bar{v}_{U_{i},m_{z}},\,s_{U_{i},m_{z}},\,\mathrm{CI}_{95}(U_{i},m_{z})\bigr)$, along with cumulative telemetry and execution metadata. This summary serves as the canonical artifact for subsequent JSON/HTML report generation (appendix E.2) and cross-proxy comparison. The caching layer eliminates repeated LLM calls across units/runs, and the plan manifest enables exact replays of component instantiation and invocation for reproducible results.
4 Metrics & Evaluation

### v2

@@ Abstract.
Large language models (LLMs) are increasingly used as human simulators, both for evaluating conversational systems and for generating fine-tuning data. However, naive “act-as-a-user” prompting often yields verbose, unrealistic utterances, motivating principled evaluation of user proxy agents. We present MirrorBench, a reproducible and extensible benchmarking framework that evaluates user proxies solely on their ability to produce human-like user utterances across diverse conversational regimes, explicitly decoupled from downstream task success. MirrorBench combines three lexical-diversity metrics (MATTR, Yule’s $K$, and HD-D) with three LLM-judge-based metrics (GTEval, Pairwise Indistinguishability, and Rubric-and-Reason), and contextualizes judge scores using Human–Human and Proxy–Proxy calibration controls. Across four public datasets, MirrorBench yields variance-aware comparisons and reveals systematic gaps between user proxies and real human users. The framework is open source [^1] https://github.com/SAP/mirrorbench and includes a command-line interface for running and managing user-proxy benchmarking experiments.
Large Language Models (LLMs), User-proxy agents, Human-likeness benchmarking, LLM-as-a-judge, Judge sensitivity, Human–LLM correlation, Lexical diversity, MATTR, HD-D, Yule’s K, Pairwise indistinguishability, GTEval, Rubric-And-Reason, ChatbotArena, ClariQ, QULAC, OASST1
copyright: acmlicensedjournalyear: 2026doi: XXXXXXX.XXXXXXXconference: Preprint Under Review; Feb 2025; isbn: 978-1-4503-XXXX-X/2026/08ccs: Computing methodologies Intelligent agentsccs: Computing methodologies Discourse, dialogue and pragmaticsccs: Computing methodologies Lexical semanticsccs: Computing methodologies Natural language generationccs: Computing methodologies Simulation evaluation
1. Introduction
@@
Deploying and improving conversational AI hinges on evaluation under realistic user interactions. Expert human-annotated utterances, while authoritative, is costly and difficult to scale for systematic evaluation. At the same time, the instruction-tuning of models for production systems demands large volumes of authentic user–assistant exchanges. To meet both needs, practitioners increasingly rely on user proxy agents, namely LLMs prompted to emulate specified human personas, to synthesize user utterances for automated testing and post-training at scale.
In practice, user proxies are now used far beyond toy role-play: they drive regression testing for assistant updates (Wang et al., 2025b), evaluate tool-use and grounding behaviors (Zhang et al., 2025), generate synthetic dialogues for post-training (Naous et al., 2025; Hathidara et al., 2025), and enable stress tests that are infeasible with human participants (Lynch et al., 2025; Zhang and Yang, 2025). However, naive “act-as-a-user” prompting often yields verbose, overly cooperative utterances that diverge from real user behavior (Zhou et al., 2024), indicating that unrealistic proxies can bias downstream evaluations and introduce distribution shift when used for data generation. This motivates a central question: How can we rigorously measure the human-likeness of a user proxy, independently of downstream task success? Here, we define a user as a human interlocutor participating in a specific conversational or task-oriented dialogue.
We introduce MirrorBench, a reproducible and extensible framework for variance-aware user-proxy benchmarking on human-likeness metrics using multiple human reference datasets across diverse conversational settings. MirrorBench evaluates proxies solely on human-likeness and deliberately separates this from task completion. We define the human-likeness of a user proxy as the extent to which its generated user utterances match real human user utterances along two complementary axes: (i) lexical similarity, operationalized via human-anchored lexical-diversity statistics (e.g., MATTR, HD-D, Yule’s $K$), and (ii) behavioral realism, operationalized via LLM-judge evaluations that assess whether the proxy’s user turns sound like a real user in context (e.g., naturalness, tone, appropriateness).
Human-likeness is inherently scenario-dependent: what sounds realistic in open-domain chat may be unrealistic in information-seeking or clarification-heavy interactions. Accordingly, MirrorBench does not aim to capture an unconstrained notion of “general” human behavior. Instead, we evaluate utterance-level realism conditioned on the same task context, asking whether a proxy produces user turns that resemble those of real users in that regime. This framing explicitly decouples user-proxy quality from assistant task success, and treats human-likeness as a measurable property of the user side of the interaction under controlled conditions.
@@
To evaluate the realism of user–proxy agents against real-world dialogues, MirrorBench packages four open-source corpora preprocessed for user-proxy evaluation: QULAC (Aliannejadi et al., 2019), ClariQ (Aliannejadi et al., 2021), OASST1 (Köpf et al., 2023), and ChatbotArena (Zheng et al., 2023). To quantify alignment in lexical diversity with human users, we implement Moving-Average Type-Token Ratio (MATTR) (Covington and McFall, 2010), Yule’s $K$ (Tanaka-Ishii and Aihara, 2015), and a Hypergeometric Distribution Diversity (HD-D) (Carmona et al., 2019). To assess behavioral realism, we employ a family of LLM-judge metrics, including GTEval (Zhu et al., 2025), Pairwise Indistinguishability (PI) (Zheng et al., 2023), and Rubric-and-Reason (RNR) (Liu et al., 2023b). Taken together, these metrics quantify the realism of the user side of the dialogue, explicitly decoupled from assistant capabilities and task success.
Our benchmark reveals that even strong LLMs exhibit systematic gaps from real users, and that these gaps vary substantially across conversational regimes. Across datasets, we observe a recurring realism-diversity tension: proxies that score highly on judge-based behavioral realism can still under-match human lexical/distributional characteristics in clarification-centric settings, motivating the use of complementary metric families. We also find that absolute judge scores and, in some cases, fine-grained orderings can shift with the judge choice, reinforcing the importance of transparent multi-judge reporting and calibration controls when comparing closely matched proxies.
In summary, our contributions are as follows: (1) a benchmark protocol for evaluating user-proxy agents for human-likeness with human-anchored normalization and calibrated judge metrics; (2) six metrics spanning lexical-diversity (MATTR, HD-D, Yule’s $K$) and judge-based realism (GTEval, Pairwise Indistinguishability, Rubric-and-Reason); (3) a unified preprocessing of four open conversational datasets for user-proxy benchmarking; and (4) an empirical study across five user-proxy LLMs, yielding actionable insights on realism-diversity tensions and judge sensitivity.
2. Related Work
User simulation and user proxies.
Simulating users to train and evaluate dialogue agents has a long history in NLP (Balog and Zhai, 2023). Early user simulators were typically goal-driven or rule-based, using predefined agendas over dialogue acts to approximate human behavior (Schatzmann and Young, 2009). More recently, LLM-based user proxies have enabled open-ended interactions that better match real conversational variability, but they also introduce failure modes such as verbosity, role drift, and inconsistent intent adherence. For instance, (Wang et al., 2025a) proposes a User Simulator with Implicit Profiles (USP) that infers latent user traits to improve realism beyond naive role-play prompting. LLM user proxies are increasingly used as evaluation drivers in interactive testbeds (e.g., tool-use and multi-turn assistant evaluation) (Lu et al., 2025; Ahmad et al., 2025; Hathidara et al., 2025), and related systems generate synthetic human–assistant conversations from seed dialogues (Zhu et al., 2025). In parallel, broader agent evaluation and simulation frameworks (e.g., AgentBench (Liu et al., 2025) and AgentScope (Pan et al., 2024)) evaluate agents in environments or support large-scale multi-agent rollouts, but do not isolate and benchmark the human-likeness of the user role as a first-class object. In contrast, MirrorBench evaluates user proxies directly by comparing their produced user utterances against real-user distributions across multiple conversational regimes, combining judge-based behavioral realism with distributional/lexical realism.
LLMs as judges.
LLM-as-a-judge evaluation prompts a strong model to rate or compare dialogue quality, reducing reliance on reference answers or large-scale human annotation (Liu et al., 2023a; Li et al., 2023). Systems such as MT-Bench/ChatbotArena (Zheng et al., 2023), G-Eval (Liu et al., 2023a), and rubric-driven evaluators such as Prometheus (Kim et al., 2024) have enabled scalable, flexible assessments by operationalizing custom criteria in natural language prompts. However, most LLM-judge setups focus on evaluating assistant outputs, whereas our setting repurposes judge-based evaluation to assess the user side and pairs it with explicit calibration controls (HH/PP) and complementary lexical/distributional metrics to better contextualize score magnitudes across datasets and model families.
3. MirrorBench
This section describes the MirrorBench evaluation methodology for evaluating user-proxy LLMs across diverse conversational datasets. Figure 1 summarizes the MirrorBench pipeline. We start from diverse conversational datasets and synthesize user goals to drive controlled, goal-conditioned rollouts between a user-proxy LLM and an assistant LLM under explicit role separation. The resulting transcripts are scored with a combined evaluation suite of judge-based behavioral realism metrics and lexical/distributional metrics, with Human–Human (HH) and Proxy–Proxy (PP) controls used to contextualize judge scores. Finally, we run robustness analyses, swapping assistants or judges and repeating runs across seeds, to quantify sensitivity and variance before reporting benchmark outputs such as leaderboards, reliability checks, and operational profiles. Detailed system architecture is described in Appendix A.
Figure 1. MirrorBench benchmark pipeline, spanning dataset preparation, synthetic rollouts, the evaluation suite, and an overlay of robustness/sensitivity analysis.
3.1. Benchmark Protocol
@@
To compare a user–proxy agent’s behavior against human references, we synthesize proxy–assistant dialogues by rolling out the user–proxy $\theta_{u}$ against an assistant model $\theta_{a}$. For each reference dialogue $d^{\mathrm{ref}}_{j}$ with user goal $g_{j}$, we condition $\theta_{u}$ on $g_{j}$ and the accumulated dialogue history to produce user turns $\hat{u}_{j,t}$, while $\theta_{a}$ produces contextual assistant replies $\hat{a}_{j,t}$. Here, we condition $\theta_{a}$ on $d^{\mathrm{ref}}_{j}$ such that the synthetic rollout trajectory follows the same path as real one. This yields a synthetic transcript
@@
3.2. Datasets & Tasks
We evaluate user proxies across four open-source conversational datasets spanning diverse domains and interaction patterns, collectively providing 795 conversations with real human user turns, enabling direct comparison between proxy-generated and real user behavior.
Table 1. Datasets used for the evaluation. All datasets are open-source and include real human user utterances.
Dataset Samples Avg Turns
ChatbotArena (Zheng et al., 2023) 195 2.5
ClariQ (Aliannejadi et al., 2021) 200 7.0
OASST1 (Köpf et al., 2023) 200 3.3
QULAC (Aliannejadi et al., 2019) 200 2.0
In Table 1, notice that the number of samples we use for benchmarking does not exhaust all the samples of the underlying open-source datasets. Instead, we perform stratified sampling over the original datasets to extract the final evaluation datasets.
For each dataset, we first normalize conversations into alternating user-assistant turn sequences, retaining only English dialogues with at least two turns. We then define dataset-specific stratification keys to ensure broad coverage: OASST1 & ChatbotArena conversations are bucketed by the number of user turns (short, medium, long), QULAC entries are grouped by topic type & facet category, and ClariQ dialogues are distributed across topic buckets & clarification-pair counts. Within each stratum, we allocate samples proportionally to the population size while enforcing a minimum-per-stratum threshold to prevent underrepresentation. Using a fixed random seed for reproducibility, we sample up to 200 samples per dataset, yielding a balanced benchmark that spans diverse conversational structures, topic domains, and interaction patterns without oversampling any single mode.
For each conversation in all four datasets, we generate a user goal description using an auxiliary LLM ($\theta_{g}$ defined in § 3.1) that summarizes the user’s intent, behavior, tone, and persona based on the real conversation. This description serves as the initialization prompt for user proxies during evaluation.
3.3. Metrics